OpenStudy (anonymous):

Can someone please check asap please Combine like terms. 7/8 x- 1/2x A.1/8 x <------------ my answer B.3/8 C.3/8 x D. x

OpenStudy (anonymous):

\[\frac{7}{8}-\frac{1}{2}\] \[\frac{7}{8}-\frac{1\times4}{2\times4}\] \[\frac{7}{8}-\frac{4}{8}\] \[\frac{3}{8}\]

OpenStudy (anonymous):

thank you so much